the wavelength(W) of a radio wave varies inversely as its frequency(F). A wave with a frequency of 200kHz has a length of 252m. What is the length of a wave with a frequency of 900kHz?
Answer by Alan3354(6092)

the wavelength(W) of a radio wave varies inversely as its frequency(F). A wave with a frequency of 200kHz has a length of 252m. What is the length of a wave with a frequency of 900kHz?
------------------
200*252 = k (a constant)
L*900 = k = 50400
L = 50400/900 = 56 meters
-----------------
-----------------
That's the answer given the data, but it's not correct.
The wavelength in meters is 300/f, f = frequency in Megahertz.
200 KHz = 0.2 MHz
300/0.2 = 1500 meters,
not 252 meters